Instrument Log

SWS                C1                  Lamont, OK Optical Trailer

Date

Technician

Comments

6/2/15

JEM

The air compressor for the dry air system has started tripping the breaker again.  The SWS is not in operation at the current time so the compressor problem is not effecting any data.  Troubleshooting determined that the compressor, regulator and start switch were not the cause of the problem and that the new motor is failing.  R/R the compressor motor and RTS the dry air system at 2000 GMT.

3/31/15

PD/KT

Mentor requests instrument and all calibration related items removed and sent so computer can be upgraded to Windows 7 and instrument calibrated.  Shut down system at 1353 GMT.  Removed *computer, **power supplies, and all LabSphere items related to calibration of the instrument.  Took to shipping department. RTS at 1533 GMT.

*Computer WD54581.

**Power Supply #1 Model LPS-200, SN 111604 18298. 

**Power supply #2 Model LPS-200, SN 120604 18417. 

3/6/15

KT

SDS reported that there was no data from the instrument for 3 hours. At 2205 GMT while trouble shooting, the computer was found to be locked up. There was no response from the mouse or keyboard. At 2210 GMT rebooted the computer. At 2215 GMT the computer came online and the program began normal operations. RTS at 2220 GMT.

1/23/15

JEM

The air compressor was found inoperable during morning rounds.  R/R the air compressor motor with a newly purchased motor.  Dry air flow returned to the SWS once the compressor was restarted.  RTS @1900 GMT.

1/23/15

KT

The Purge Air Pressure was reading 0. The main outside Air Compressor was off. At 1512 GMT checked the breaker. Breaker didn't appear to be tripped but reset the breaker. No change in status, the Air Compressor did not come on. Pressed the reset switch on the compressor and reset the breaker again. No response from the Air Compressor. Trouble shooting in progress. RTS at 1530 GMT.

1/12/15

JEM

Ken Teske notified me that on morning rounds he found the air compressor inoperable again.  Troubleshooting found the start capacitor had failed again.  Due to continued issues with the air compressor motor, I ordered 2 new motors, one for use and one for a spare.  Was able to find another capacitor that was close enough to the correct size to get the motor running again.  Will install the new motor either when it arrives or when the current motor fails again, whichever happens last.  RTS @2000 GMT.

1/12/15

KT

The main air compressor that provides air for the Dry Air system was not operating. During inspection the breaker for the main air compressor was tripped. Attempted to reset the breaker twice and still could not get the air compressor to return to service. Trouble shooting is in progress. RTS at 1717 GMT.

12/9/14

JEM

The air compressor stopped working again approximately 1 hour after it was returned to service on 12/8.  Tried resetting the breaker and again it would trip after a few seconds of operation.  Disassembled the motor and checked the internal centrifugal start switch again and it was not stuck and was working properly.  Suspect a problem with the motor.  R/R the motor with the spare.  Reset the breaker after installing and connecting the spare motor.  The breaker would immediately trip and the motor would not even try to run.  Removed the motor and took both motors to the repair lab for further inspection.  The spare motor would also trip the breaker in the repair lab and further inspection found a burnt spot in one of the windings causing a short.  Powered on the original motor in the lab and it ran just fine and current draw was only 7 amps.  Returned the original motor to the optical trailer air compressor and again it would trip the breaker after a few seconds of operation.  Current draw was measured at ~40 amps.  Inspection of the cut in/out switch connected to the high pressure cutoff revealed that all of the contact pads had a large amount of carbon buildup and damage from arcing.  R/R the cut in/out switch with a new one.  Due to all of the times the breaker has tripped in the past the breaker was also replaced with a new one.  Powered the compressor on at 2200 and it ran normally.  Dry air flow returned to the SWS.  RTS @2200 GMT.

12/8/14

JEM

The compressor for the dry air system has failed again.  Tried resetting the breaker several times and the breaker would trip after a few seconds of operation.  Disassembled the air compressor motor and the internal centrifugal switch was stuck in the on position, not allowing the start winding to disengage once the motor was up to speed causing excessive current draw.  Replaced the internal centrifugal start switch with a new one.  Reassembled the motor and powered on the compressor.  Dry air flow returned to the SWS.  RTS @1830 GMT.

12/1/14

JM/CM

The dry air compressor has tripped the breaker several times in the past few weeks.  Replaced the start capacitor for the compressor motor at 1900 GMT.  One of the terminals broke off the original capacitor when it was disconnected, so it may have been the source of the problem.

12/1/14

KT

The Outside Air Compressor was not running. At 1520 GMT found that breaker had been tripped. Reset the breaker and the Air Compressor powered back on. RTS at 1535 GMT.

11/12/14

JEM

Was notified by site ops that the air compressor used to supply slow to the SWS is not working.  The circuit breaker was found tripped and when trying to reset the breaker it trips again after about 1 second.  Troubleshooting found that the start capacitor on the motor has failed.  R/R the start capacitor and reset the circuit breaker.  Compressor began working normally.  After the compressor pressured up, observed the flow return to normal.  RTS @1830 GMT.

10/1/14

CAM

The dry air system compressor motor failed on 9/30.  Removed the failed motor and replaced with a spare rebuilt motor.  Removed smaller compressor that had been temporarily plumbed in yesterday and placed the original compressor back in service at 2000 GMT.

9/30/14

CAM

Technician found during daily rounds that the compressor motor for the dry air system had failed.  Temporarily connected a smaller compressor to the dry air system until the large compressor can be repaired.  RTS at 2115 GMT.

8/14/14

CAM

The RH inside the SAS-He refrigerator has risen over the past several weeks and is at times near 60%.  Found that SAS-He, SAS-Ze, SORTI, and SWS were all on the same regulated dry air line.  There was another regulated air line that had only RSS on it.  Reconfigured the air lines so that SAS-He and SORT were on a line by themselves.  RH dropped to near 0% for both SAS instruments and appears to be holding.  RTS at 2130 GMT.

5/28/14

Mentor

Instrument mentor on site and reinstalled SWS.

5/6/14

CAM

Mentor requests SWS removed and sent for calibration.  Shut instrument down at 1830 GMT and removed collector.  Fiber was left connected to collector and computer/chassis during removal as per Mentor's request.  Collector, computer/chassis, fiber and sun screen were taken to shipping dept.  No SN or model numbers are listed.  WD 54581. 

3/18/14

KT

The sensor head housing needed to be rotated 180 degrees. At 1449 GMT rotated the sensor head housing 180 degrees. System is online and operational. RTS at 1504 GMT.

2/12/14

KT

The Purge Air flow was reading 0. While troubleshooting it was discovered that the filter was bad and needed to be changed out. Removed and replaced filter. The flow value returned to 18. RTS at 1915 GMT.

1/30/14

CAM

Kay Electric needed to shut the power off briefly while they worked on the meter.  Power was shut off at 1614 GMT and restored at 1623 GMT.  The optical trailer UPS died during that time, causing instruments to lose power.  Power cycled UPS at 1647 GMT to restore power.  SWS restarted and RTS at 1649 GMT. 

5/8/13

CAM/JEM

Instrument was removed on 4/17 so the wooden decking and stairs could be replaced with metal.  Construction work is complete and instrument needs to be reinstalled.  Installed SWS and powered on.  RTS at 1524 GMT.

4/17/13

CAM/JEM

The optical trailer wooden decking and stairs are to be removed and replaced with metal decking and stairs.  Need to remove all instruments so removal of old decking can start tomorrow. Stopped SWS program and powered down at 1400 GMT.  Removed the sunshade and instrument and stored inside trailer.  Covered roof opening with plastic.  Instrument will remain offline until new deck and stairs are complete. 

3/19/13

CAM/KT

SWS returned from being calibrated and needed to be installed.  Installed SWS and powered on at 2040 GMT.  There was no video displayed on the monitor.  Removed chassis cover and found video connection was loose inside.  Reseated cable and video signal returned.  Replaced cover and RTS at 2045 GMT.  WD54581

3/6/13

CAM/PAD

Mentor requests instrument removed and sent for calibration.  Shut down system and removed collector.  Fiber was left connected to collector and computer/controller during removal as per Mentor's request.  Collector, computer/controller, fiber, and sun screen were taken to shipping dept.  No SN or model numbers are listed.  WD 54581.  RTS at 2254 GMT.

2/26/13

CAM

The site lost power at approximately 0350 GMT on 2/26 due to a severe winter storm.  Powered optical cluster UPS off at 1730 GMT.  Power was restored at 1920 GMT.  Restored UPS power at 1945 GMT. Instrument RTS status O.K and data collections resumed once SDS operations were restored.

1/14/13

CAM

After reinstalling the SWS, noticed the dry air system was not providing air.  T/S found the copper air line between the compressor and the building was plugged with ice.  Disconnected copper line at building entrance, thawed the ice, and cleared water from line.  Reconnected line and confirmed air flow.  RTS at 2030 GMT.   SAS-He, RSS and SORTI were also affected by lack of air.

1/14/13

KT

The SWS instrument was back from calibrations and needed to be installed. At 1900 GMT installed the SWS instrument. Powered system on at 1940 GMT.

12/3/12

KT

Instrument mentor requested that the instrument be removed and shipped out for calibrations. At 2005 GMT stopped the program and powered down the SWS instrument. Removed the SWS instrument and took down to shipping and receiving to be packed and shipped out. RTS at 2030 GMT. WD# 54581

11/20/12

CAM

Local electrical service was shut off at 1410 GMT so that Kay Electric could upgrade its power lines from 7.2 KV to 14.4 KV in Grant County.  SWS was gracefully shut down at 1349 GMT in anticipation of power outage.  Power was restored at 2115 GMT after electrical upgrades were complete.  Restarted SWS and RTS at 2149 GMT.

11/13/12

CAM

Site power was partially shut down at 1400 GMT in order to replace the 7 transformers that service the buildings and all instruments south of the main compound.  Power was restored at approximately 1930 GMT after all transformers were replaced.  SWS auto-started status O.K.

10/30/12

KT

The SWS computer was locked up.  At 1540 GMT power cycled the computer. RTS at 1555 GMT. Instrument computer online and operational.

6/20/12

PAD

Found program locked with elapsed time of 7,044 shown on program clock.  Computer is not responsive to keyboard or mouse.  Power cycled computer and program restarted correctly.  RTS at 1517 GMT.

6/7/12

KT

Last SDS collection was at 1700 GMT.  At 2052 GMT, found computer program locked.  Computer is not responsive to keyboard or mouse.  Power cycled computer at 2052 GMT.  Program started correctly after computer powered up.  RTS at 2107 GMT.

5/17/12

PAD

At 1754 GMT, observed program locked at 1191 elapsed time.  Powercycled computer.  Program started correctly.  RTS at 1809 GMT.

4/25/12

PAD

Program had locked at elapsed time 788 shown.  Shut down and restarted computer.  Program started correctly.  RTS at 1351 GMT.

4/16/12

PAD

SWS program is running but Si graph shows no signal, flatlined at zero.  Error is shown:  Error reading Si Spectrometer! error:  DACQERROR_TIMEOUT_EOS_Scan pixel: 0.  Cleared error but Si signal remained at zero.  Exited program and restarted computer.  Program started correctly.  RTS at 1544 GMT.

4/3/12

PAD

Last SDS collection was at 1600 GMT.  Program is locked and computer is not responsive to keyboard or mouse.  Power cycled computer and program started correctly.  RTS at 2050 GMT.

3/22/12

CAM

No data had been collected from SWS since 1500 GMT.  Found the program was locked up and computer was unresponsive to mouse or keyboard.  Elapsed time was stuck at 9242.  Power cycled computer at 2128 GMT.  Program restarted correctly and RTS at 2129 GMT.

3/5/12

PAD

Last SDS collection was at 1700 GMT.  Found program had locked and computer is not responsive to keyboard or mouse.  Powercycled computer.  Program started correctly.  RTS at 2232 GMT.

2/23/12

PAD

Program had locked after morning startup after 2,920 sec elapsed time.  Computer is not responsive to keyboard or mouse.  Powercycled computer.  Program started correctly.  RTS at 1719 GMT.

2/6/12

PAD/KT

System has been returned to CF and mentor requests system installed.  Installed Labsphere plate, sphere, lamp power supply, sphere controller, and computer/controller.  Plate model US-120-SF, SN 03290518817.  Lamp power supply model LPS-200, SN 111604 18298.  Sphere controller model SC-5500, SN 012405 18605.  Computer/Controller WD54581.  Started system and set integration times as directed by mentor at 240 and 100 ms.  RTS at 1920 GMT.

11/4/11

C. Flynn

Mentor pulled instrument to be shipped out for repairs.

11/2/11

C. Flynn

Mentor on site.

9/26/11

PAD

Purge Gas Generator located in Optical trailer has been repaired.  Unit needs to be RTS.  Installed Whatman FT-IR Model 75-45-12VDC, SN 7545-0236 Purge Gas Generator and connected inlet line.  Started unit and opened inlet valve to allow air to flow through system.  Vendor recommends unit allow to run 24 hours before connecting outlet line.  Will connect outlet line tomorrow.  Instruments affected are SWS, RSS, and SORTI.  RTS at 2120 GMT.

9/8/11

PAD

Error shown:  Error reading Zenith Si Spectrometer!  error DACQERROR_TIMEOUT_EOS_SCAN .  Cleared error and rebooted computer.  RTS at 1540 GMT. 

8/4/2011

PAD

Purge Gas Generator located in Optical trailer that services RSS, SWS, and SORTI is not operating correctly.  Condensation is getting past unit and into lines going to instruments.  Unit needs to be removed and sent for repairs.  At 1520 GMT, shut down compressor and removed unit.  Whatman model FR-IR 75-45-12VDC, SN 7545-0236.  RTS at 1600 GMT.

7/29/11

SD

Airflow gage had water inside of it also both air valves condensation site windows were full of water.  Purged airflow system

6/24/2011

JEM

Program is locked.  Error shown:  Error writing Spectrometer data to file.  Also in the upper left corner of the screen a linux message is showing through the program window and it states "journal commit I/O error".  Exited the program from the menu and noticed while the computer was shutting down a message scrolled by that said something to the affect of hda0 or hda1 has bad sectors and also there was an error unmounting the drive.  Powered the computer back on and the program came up normally.  RTS @1439 GMT.

6/23/11

PAD

Program is locked.  Error shown:  Error writing Spectrometer data to file.  Cleared error message and powercycled computer.  Program started correctly.  RTS at 1348 GMT.

6/21/11

PAD

Program is locked and error shown:  Error writing Spectrometer data to file.  Computer was not responsive.  Power cycled computer and program restarted correctly.  This is an ongoing issue.  Troubleshooting with mentor at this time.  RTS at 1446 GMT.

6/20/11

PAD

At 1255 GMT, SDS shows no SWS collections for 23 hours.  At 1734 GMT, error was shown:  Error writing Spectrometer data to file.  Cleared error with space bar but program is locked.  Power cycled computer and program started correctly but at 1956 GMT, noticed same error message.  Notified mentor.  Troubleshooting at this time.

6/16/11

PAD

Computer program is locked up and not responsive to keyboard or mouse.  Error shown:  Error writing Spectrometer data to file.  Shut down/ restarted computer.  Program started correctly.  RTS at 1533 GMT.

6/15/11

PAD

Computer/controller was removed on 5/5/11 and sent for repairs.  It has been returned and needs to be installed.  Installed computer/controller and connected to instrument at 1514 GMT.  RTS at 1529 GMT.

5/5/11

PAD

SWS InGaAs temperature is -24.5C and should be -10C.  Previous troubleshooting has failed to identify problem.  Mentor requests computer/spectrometer removed and sent for repairs.  Removed computer/spectrometer, WD54581 and took to shipping department.  RTS at 1523 GMT.

5/4/11

PAD

Program has stopped and error is shown:  Error writing Spectrometer data to file.  Exited program, shut down and restarted computer.  RTS at 1422 GMT.

5/2/11

PAD

InGaAs temperature is -24.5C.  Mentor requests computer controller box opened and inspected for possible causes of temperature being out of range.  Shut down computer, disconnected power, and removed cover.  Did not find any obvious reasons for temperature reading.  Reseated ribbon cable between InGaAs controller board and TEC cooler.  Started up instrument and temperature returned to -24.5C reading within 10 minutes.  Notified instrument mentor.  RTS at 1905 GMT.

4/8/11

PAD

Found program locked.  Kernel panic- not syncing error is shown.  Computer is unresponsive to keyboard or mouse.  Power cycled computer, program started and show a data directory error.  Shut down/ restarted computer and program started correctly.  RTS at 1547 GMT.

4/7/11

PAD

Found program locked.  Kernel panic- not syncing error is shown.  Computer is unresponsive to keyboard or mouse.  Power cycled computer, program started and show a data directory error.  Shut down/ restarted computer and program started correctly.  RTS at 1337 GMT.

3/31/11

PAD

Program had locked.  Kernel panic- not syncing error is shown.  Computer is not responsive to keyboard or mouse.  Power cycled computer and restarted program.  RTS at 1439 GMT.

3/30/11

PAD

Program had locked.  Kernel panic- not syncing error is shown.  Computer is not responsive to keyboard or mouse.  Power cycled computer and restarted program.  RTS at 1521 GMT.

3/28/11

PAD

Program had locked.  Kernel panic- not syncing error is shown.  Computer is not responsive to keyboard or mouse.  Power cycled computer and restarted program.  RTS at 1417 GMT.

3/24/11

PAD

Mentor requests ribbon cable between TEC cooler and controller R/R.  Computer was shut down at 1503 GMT due to locked program.  R/R ribbon cable as instructed by mentor.  Powered up system at 1528 GMT.

3/24/11

PAD

Program locked on 3/23/11 at 1749 GMT.  Error shown:  Kernel panic- not syncing: drivers/ide/ide-io.c:1541: spin_lock (drivers/ide/ide.c:c03dafe8) already locked by drivers/ide/ide-lib.c/454. (Not tainted).  Computer is not responsive to keyboard or mouse.  Shut down computer and restarted at 1528 GMT.

3/15/11

CM/MK

SWS was shipped back to the site after being repaired and needed to be installed.  Also included was a new aluminum ring flange to replace the old white nylon ring between the light collector and roof port.  The holes in the new ring did not line up with the holes in the roof port.  Drilled new holes in port to match the ring.  The gasket supplied with the ring was too thin to adequately seal the slightly uneven gap between the port and ring.  Cut a new gasket from thicker bulk material for a tight seal.  Installed SWS, powered on and RTS at 1510 GMT. 

11/17/10

CAM

Instrument mentor requested the SWS be removed and returned to PNNL to have the InGaAs spectrometer installed and the instrument calibrated.  Stopped the program and powered off at 1521 GMT.  Removed instrument to shipping to be returned to mentor.  Returning the computer/spectrometer chassis, light collector and baffled dome.  WD54581

11/10/10

CAM

AC power will be interrupted for construction near the main compound.  Need to gracefully shut the instrument down prior to power outage.  Powered instrument down at 2152 GMT.  Restarted instrument after AC power was restored.  RTS at 2327 GMT.

10/28/10

PAD

Program had locked on 10/27/10 at 1934 GMT.  Error message across top of monitor screen:  Kernel-panic.  Not synching.  Powercycled computer and started program.  RTS at 1421 GMT.

10/26/10

PAD

SWS program had locked up 54 seconds after starting.  Computer is not responsive to keyboard or mouse.  Powercycled computer and restarted program.  RTS at 1451 GMT.

9/9/10

PAD

SWS program has locked up.  Program show 1701 sec on elapsed time clock after program started today.  Computer is not responsive to keyboard or mouse.  Power cycled computer.  Instrument started taking measurements.  RTS at 1424 GMT.

7/23/10

Flynn

The instrument computer is inoperative.  Troubleshooting found that the drive installed was not responding. Replaced the drive with the original drive and restarted at 1700 GMT. Worked with the SDS team to correct coms settings and data started transferring at 1730 GMT. The instrument is still only operating on one channel.

6/21/10

PAD

Instrument mentor requests operating and data cards reseated in computer.  Shut system down at 1515 and reseated cards.  Restarted instrument and notified mentor.  RTS at 1530 GMT.

6/1/10

PAD

At 1458 GMT error was shown:  Error creating Spectrometer Data File.  Rebooted computer but same error is shown.  Normally error will clear after program is exited and computer rebooted.  Contacted mentor.  At 2107, mentor requested data directory changed to /media/idedisk/sciflt/.  Working with mentor, deleted files to open hard drive space and changed directory.  RTS at 2147 GMT.

4/26/10

PAD

InGaAs temperature is 27.1 C and out of normal operating range.  Shut system down and notified instrument mentor.  Troubleshooting at this time.  RTS at 1441 GMT.

4/19/10

PAD

Last SDS collection was at 1701 GMT.  Found computer program locked and unresponsive to keyboard or mouse.  Powercycled computer and restarted program.  RTS at 2247 GMT.

2/23/10

PAD

Window in SWS sensor enclosure has a film on inside.  Window needs to be removed and cleaned.  Removed and cleaned window at 1934 GMT.  RTS at 1949 GMT.

2/10/10

PAD

Computer program is locked and unresponsive to keyboard or mouse.   Power cycled computer and restarted program.  RTS at 1529 GMT.

1/31/10

PAD

Last SDS collection was on 1/28/10 at 2201 GMT.  Computer had locked at 2231 GMT.  Computer is not responsive to mouse or keyboard.  Power cycled computer and restarted program.  Cleared ice/snow from tube above instrument window.  RTS at 2046 GMT.

10/30/09

MK

The SWS computer is locked up. Power cycled the computer at 1423 GMT. Error finding data directory. Shut down and restarted again at 1505 GMT. Data directory problem corrected itself. RTS at 1510 GMT.

9/17/09

PAD

Program is locked.  Computer is unresponsive to keyboard or mouse.  Power cycled computer.   Program started taking measurements.  RTS at 1535 GMT. 

9/15/09

PAD

Mentor requests Zenith InGaAs integration time changed from 230 to 220ms.  Made changes as requested at 1530 GMT. 

9/14/09

PAD

Program is locked.  Computer is unresponsive to keyboard or mouse.  Found computer is connected to raw power.  Connected computer to UPS outlet and restarted.  Program started taking measurements.  RTS at 1525 GMT.

4/8/09

PAD

Entered Optical building and heard air leak from dry air system for SWS and Sorti.  Found leak was o-ring on drain valve of inline filter.  Replaced o-ring and leak was stopped.  Found another leak in tubing between regulator and filter.  R/R with copper tubing.  RTS at 1545 GMT.

2/3/09

PAD

Mentor has determined instrument has been measuring scattered light affecting data quality.  SWS sensor head was removed and sent to mentor on 1/29/09.  Mentor requests computer, calibration sphere, controller, power supply, and all necessary bulbs, cables, etc., sent to him.  Removed components as instructed at 1922 GMT and took to shipping dept.  RTS at 2032 GMT.

12/9/08

PAD

Program stopped after 1 second.  CF was without electricity between 1014 and 1222 GMT.  Exited program and restarted computer.  Program restarted correctly.  RTS at 1612 GMT.

12/2/08

PAD

Instrument was returned to CF after calibration at NASA-Ames.  Mentor requests instrument installed and calibrations run on site.  At 1830 GMT, Chris Martin and I installed instrument and performed calibrations as instructed.  After calibrations, instrument was RTS at 2140 GMT.

11/12/08

CM/PD

Instrument mentor requested the SWS be removed and sent to NASA Ames for calibration.  Shut the SWS down at 1916 GMT.  Removed the SWS, computer, integrating sphere, calibration bench, sphere controller and lamp C power supply for shipment to NASA. Also returning the fiber optic cable, alignment HeNe laser, photo diode and cable, 4 lamps, 1 lamp power cable and 3 power cords.  System will remain offline pending return of all components.

10/15/08

MK

InGaAs temperature on SWS has been fluctuating.  Mentor requested connections checked for corrosion and reseated between InGaAs temperature controller and TEC.  Checked connections and found no corrosion.  Thermistor between pins 5 and 6 on cable between controller and TEC  measured 5.96 K, and TEC measured 10.3 Ohms.  Powered up instrument and computer  at 1438 GMT and InGaAs temperature stabilized at -10.1 C.  Will continue to monitor.  RTS at 1504 GMT.

8/25/08

PAD

SWS InGaAs temperature is 24.4C.  Fluctuating InGaAs temperature has been an issue for some time.  Previous CMs of replacing InGaAs controller and power supply has not fixed issue.  At 1517 GMT, stopped program and shut down/ restarted computer/ controller.  After startup, InGaAs temperature was still at 24C.  Shut down instrument and notified mentor.  RTS at 1532 GMT.

8/20/08

PAD

InGaAs temperature readings have been fluctuating out of range for some time.  Mentor request temperature controller power supply R/R.  R/R Lambda part # LWD30-0512 AC to DC power supply.   InGaAs temperature stabilized at -10.3C with no out of range fluctuations observed.  Will continue to monitor.  RTS at 1908 GMT.

8/18/08

PAD

Program locked 7,058 seconds after starting.  Program clock is the only thing updating on monitor screen.  Computer is not responsive to mouse or keyboard.  Power cycled computer and program started.  Program came up with a Media Disk File error and I had to exit/ restart program.  RTS at 1424 GMT.

7/14/08

MK

CF lost power on 7/13/08 between 0028 and 0137 GMT. Instrument locked up. Restarted computer at 1512 GMT.

7/11/08

PAD

Found program locked up 14,158 sec after starting.  Date and time are correct.  Computer is not responsive to keyboard or mouse.  Power cycled computer.  RTS at 1513 GMT.

7/10/08

PAD

SWS program locked up 12,922 seconds after starting today.  Date and time are correct but no other values are updating.  Computer is not responsive to keyboard or mouse.  Previous  lock-ups found mentor is able to remotely access computer when it is not locally accessible.  Notified mentor as requested.  Later check of DSView indicates mentor was successful in rebooting computer.   RTS at 1457 GMT.

7/3/08

PAD

Program locked up 113 seconds after starting today.  Computer is not responsive to keyboard or mouse.  Power cycled computer and program restarted correctly.  RTS at 1441 GMT.

7/2/08

PAD

InGaAs temperature is periodically spiking.  Troubleshooting with mentor has determined InGaAs temperature controller is faulty and needs to be replaced.  At 1359 GMT stopped program, shut down computer, and disconnected power.  Removed controller, part # 11-0106550-02, No SN, and installed controller, part # 11-0106550-02, SN 734.  Checking on instrument at 1430 shows temperature spiked twice within the last 30 min.  Will continue to monitor InGaAs temperature. RTS at 1419 GMT.

6/16/08

PAD

CF experienced a power outage at 1834 GMT.  After power was restored, SWS program locked on restart.  At 2011 GMT, stopped program and rebooted computer.  Program restarted correctly.  RTS at 2026 GMT.

5/12/08

PAD

Notified by mentor Spectrometer InGaAs temperature is fluctuating.  Mentor requested InGaAs cooler fan checked to make sure it is running.  At 1442 GMT, checked InGaAs temperature graph.  No spikes detected in temperature for last hour.  Removed SWS controller cover and could hear a clicking sound synchronized with a blinking light on the InGaAs controller.  Temperature was spiking high and low.  External power supply previously installed to troubleshoot shows 5.002 V but is cycling between .6 and 1.1 amp.  Checked wires at controller but found no issues with connections.  Stopped fan on InGaAs cooler for 5 sec with no temperature spikes observed.  Notified instrument mentor.  RTS at 1503 GMT.

4/28/08

PAD

Notified by mentor InGaAs temperature is -24C.  External power supply is currently connected supplying power to InGaAs temperature controller.  CF lost power for approximately 4 hours on 4/26/08 at about 0000 GMT.  Found instrument had restarted but external power supply had power but 0V output.  Reset power supply to 5.003V output.   Notified instrument mentor.  RTS at 1339 GMT.

4/18/08

PAD

Spectrometer InGaAs temperature has been spiking.  Mentor request external power supply connected to temperature controller to troubleshoot possible power supply failure.  At 1711 GMT, Mark Klassen and I connected external power supply as instructed.  Powered up and restarted SWS at 1746 GMT.  After startup, observed TEC cooler fan on InGaAs spectrometer was not turning.  Light touch was enough to start fan.  Notified mentor and will leave power supply connected until Monday.  RTS at 1811 GMT.

4/1/08

PAD

Air compressor that supplies drying air to SWS and Sorti instruments located in Optical trailer shows 0 psi.  Reset breaker but compressor did not restart.  Notified facilities. 

3/12/08

PAD

No air pressure at inline dryer supplying SORTI and SWS instruments.  Reset breaker for air compressor in optical trailer.  Air compressor started and pressure stabilized.  RTS at 1543 GMT.

1/2/08

PAD

Instrument has been returned to CF after calibration and needs to be installed and returned to service.  At 1845 GMT, installed instrument, computer, power supply and all related calibration equipment.  Started program and notified instrument mentor.  RTS at 1945 GMT.

10/26/07

PAD

Program has locked up.  Elapsed time shows 12,215 seconds.  Computer is not responsive to mouse or keyboard.  Cannot open virtual window.  Contacted instrument mentor who remotely accessed and rebooted computer.   RTS at 1529 GMT.

10/5/07

PAD

Si temperature controller is not working correctly and needs to be replaced.  Stopped program and shut down computer.  R/R Si temperature controller located inside computer/ controller box.  Si temperature settled around 27.2 C.  RTS at 1432 GMT.

9/27/07

PAD

Si temp is low and registering approximately room temperature.  Mentor requests T/S to verify heater controller has failed.  Assisted Mark Klassen in T/S as per instructions.  T/S found heater controller had failed.  Notified instrument mentor.  RTS at 1938 GMT.

7/31/07

MK

The program was locked up. Shutdown computer, restarted with error stating problem with creating spect data files.  Contacted Mentor. Shutdown system again and restarted at 1720 GMT. System returned to normal operation.

7/30/07

PAD

Program found locked at 257 seconds elapsed time.  Program clock is only value updating.  Status shows recording dark signal.  Unable to access virtual window to shut down computer.  Computer is not responsive to keyboard or mouse.  Left voicemail with instrument mentor.  At 1513 GMT, power cycled computer.  After program restarted error message was displayed:  Error creating spectrometer data file!  Please check directory settings.  Exited program and shut down computer.  Computer and program restarted correctly.  Notified instrument mentor.  RTS at 1528 GMT.

7/17/07

PAD

Two error messages are displayed:  Quitting will abort all running Vi's and  Bruker FTIR.  AQP Read Timeout.   Canceled error messages but display keeps showing Bruker FTIR error message.  At 1459 GMT stopped program and shut down computer and instrument.  Powered up instrument and then computer.  Error message came up but did not return after canceling.  RTS at 1514 GMT.

7/12/07

PAD

Error message displayed:  Error creating Spectrometer data file!  Please check directory settings.  Only thing active and updating on program is clock.  Storms in area this morning and Optical trailer UPS was in alarm.  Exited program and restarted computer.  Computer and program restarted correctly.  RTS at 1547 GMT.

6/22/07

PAD

At 1922 GMT found SWS program was locked.  Only thing updating is the clock.  Cannot open virtual window.  Notified instrument mentor.  Mentor remotely accessed computer and rebooted.  Program restarted correctly.  RTS at 1942 GMT.

6/14/07

PAD

Inlet pressure on dryer controller for SWS and SORTI shows 0 psi.  Reset breaker for air compressor.  RTS at 1445 GMT.

6/13/07

PAD

Inlet pressure on dryer controller for SWS and SORTI shows 60 psi.  Reset breaker for air compressor.  RTS at 1508 GMT.

6/1/07

MK

The SWS computer is not running. Instrument mentor rebooted remotely at 1522 GMT.

5/29/07

MK

The power at the CF had failed some time over the weekend causing the UPS to shutdown due to low voltage on batteries. Powered on UPS at 1434 GMT, restarted SWS computer and software at 1440 GMT.

4/17/07

PAD

Si Temp is reading 21.1 C. and is normally 27.0 C.  Called instrument mentor who requested program stopped and computer/ controller power cycled.  Shut down program and power cycled computer/ controller at 1422 GMT.  No change in Si Temp.  Troubleshooting at this time.  RTS at 1432 GMT.

4/2/07

PAD

Computer has power but is not responsive to keyboard or mouse.  Removed computer from raw power and plugged into UPS power.  Removed monitor from UPS power and plugged into raw power.  Restarted program.

3/20/07

PAD

Storms at CF tripped UPS located in Optical trailer.  Last successful collection was at 0300 GMT.  Following instruments were affected:  AERI01, MWR, RSS, SWS, TSI-880, SORTI01, and VCEIL25K.  Facilities reset UPS at 1600 GMT.  Restarted computers that did not restart after power was restored.  RTS at 1635 GMT.

1/24/07

PAD

Computer program has locked up.  Program clock is the only thing updating on monitor screen.  Computer is not responsive to mouse or keyboard.  Power cycled computer at 1713 GMT.  Program restarted correctly.  RTS at 1728 GMT.

1/18/07

PAD

Program has locked up.  Clock is the only thing updating.  Computer is not responsive to mouse or keyboard.  Cannot open virtual window.  Power cycled computer at 1622 GMT.  Program restarted correctly.  RTS at 1637 GMT.

1/11/07

PAD

Storm moving into area with freezing rain predicted.  Dan Rusk request protective cap placed on lens.  Put protective cap on at 2225 GMT.  RTS at 2240 GMT.

10/30/06

CAM         

BCR-01288 requests that the SWS be returned to NASA-Ames Research Center to recalibrate the instrument with the Airborne Science Facility's 30 inch integrating sphere, test different integration times to eliminate a low signal to noise ratio, and to do a stray light test.  Shut the system down at 1947 GMT.  Removed the SWS light collector and electronics box, integrating sphere's lamp C *power supply and the integrating sphere **controller.  The integrating sphere and its mounting bench had been returned to Ames previously.  Also included in the shipment will be the fiber optic cable, photo diode and associated BNC cable, 4 lamps, one lamp power cable, 3 AC power cords.  Covered the port hole with a blank plate as well as a plastic bag.  System is offline at this time.  *Old - 120604 18418  **Old - 012405 18605

10/10/06

PAD

Last successful collection was on 10/9/06 at 1402.  Computer had locked at 1413 GMT.  Computer is not responsive.  Error message displayed reads:  Kernel panic- not syncing:  drivers/ ide/ ide- io.c.1541: spin- lock (drivers/ ide/ die.c: c03dafe8)  Already locked by drivers/ ide/ ide- lib.c/ 454.  Not tainted.  Power cycled computer and tried to restart program.  Error message displayed:  Error creating spectrometer data file!  Please check directory settings.  Called instrument mentor.  At 1907 GMT, shut down computer and swapped data card with spare.  System restarted correctly.  Notified instrument mentor.  RTS at 1922 GMT.

9/28/06

PAD

At 1623 GMT, observed SWS display was not updating.  Program had locked.  Computer time was correct.  Tried to open a virtual window to reboot computer but it would not open.  Mentor was able to remotely access computer and reboot.  SWS program restarted correctly.  RTS at 1638 GMT.

8/31/06

CAM

Instrument mentor requested the SWS calibration bench, integrating sphere and HeNe laser be returned to NASA Ames for modifications.  Removed all the required components and boxed them for shipment.  Delivered to S/R department.

8/23/06

PAD

Warren Gore requests plate surrounding the SWS dome changed out with one sent by him.  Mark Klassen and I R/R with plate provided at 1747 GMT.  RTS at 1802 GMT.

8/2/06

PAD

At 1445 GMT, observed computer program was locked and computer was not responsive.  Program had started and locked after elapsed time of 162 seconds. Was unable to open virtual window to reboot computer.  Power cycled computer but program did not restart correctly.  Error displayed:  Error Creating Spectrometer data file!  Please check directory settings.  Opened virtual window and rebooted computer as per instructions.  Program restarted correctly.  Notified instrument mentor.  RTS at 1500 GMT.

7/21/06

PAD

Started lamp C power supply for calibration at 1832 GMT.  Finished calibration at 1855 GMT.  RTS at 1902 GMT. 

6/28/06

PAD

Air conditioner is not working in optical trailer.  Notified instrument mentor who requests instrument shut down if si temp reaches 30 C.  At 1531 GMT, si temp was 30.9 C.   Shut down computer and instrument at 1532 GMT.  Placed protective cap over dome.  Notified instrument mentor.  RTS at 1547 GMT. 

4/26/06 -4/28/06

Mentor

Instrument mentor on site installing the new SWS system.  Site Ops received training and instrument operational on 4/28.

11/21/03

CAM

In light of the recent failure of the SWS, BCR-00763 states: Following a discussion of ARM's needs for a shortwave spectrometer at the recent ARM Coordination Team meeting, the consensus decision was to remove the SWS from service and to proceed with field trials of the new shortwave spectrometer Peter Pilewskie is helping develop as soon as that instrument is ready.  Removed *SWS instrument, computer and dome and placed in storage.  Notified mentor of the actions taken.  Also returned to Margaret Skokan the laptop computer/power supply that was on loan from ASD.  *Old - 682

10/30/03

CAM

Received a laptop computer from ASD, which contains software to help diagnose the problem of the instrument not communicating with computer.  Connected the SWS instrument to the laptop at 2010 GMT.  Tested the instrument as per phone instructions from Margaret Skokan.  The instrument still would not communicate with the laptop.   Suspect an instrument hardware failure.  Margaret will discuss the matter further with ASD Engineers and get back with us about possibly returning the instrument for repairs.  Instrument remains down at this time.

10/29/03

CAM

The SWS program has locked up again.  Attempts to restart always results in the program locking up after about 4 seconds with no response from instrument.  Margaret Skokan from ASD recommended installing the Fr software on a different computer and see if the instrument will respond.  Loaded Fr software on spare Windows 98 laptop and tested with same results.   Margaret will ship a computer from ASD with additional diagnostics software.  System remains down at this time.

10/27/03

PAD

SWS last successful collection was on 10/27/03 at 0002 GMT.  Computer was running and responsive.  Computer program had locked up at 0001 GMT.  Restarted program at 1500 GMT.  Program locked up 4 seconds later.  Restarted program again and it ran 4 sec.  Called John Pommier to troubleshoot problem.  Rebooted computer and power cycled instrument at 1531 GMT.  Program locked up after 4 sec.  Removed computer panel and checked that cards were correctly seated.  Program locked up.  Restarted program at 1950 GMT. Computer program is currently running.  Troubleshooting with mentor at this time.

10/15/03

MR

The instrument mentor wants the desiccant changed out of the dome and a new O-ring put in order to see if this will cure the problem with the moisture in the dome.  Changed the desiccant and put in a new O-ring.  RTS at 2000 GMT.

10/9/03

PAD

Condensation inside dome.  Removed dome and cleaned condensation.  RTS at 1530 GMT.

10/6/03

PAD

SWS dome has been having condensation inside dome when humidity is high.  Replaced desiccant at 1725 GMT.  RTS at 1740 GMT.

10/2/03

CM

The SWS Fr_Auto program was found locked up.  This has recently become common.  The computer itself does not lock up, just the program.   Suspect memory or data transfer problem.  Found that there were 1000 1-minute data files located in C:\ Fr\data.  These files were dated 10/29-10/30/01.  Moved these files to temporary folder C:\Chris\data.  Restarted computer and RTS at 1510 GMT.  Monitored instrument status throughout the day and it remained operational.  It is not certain why, or if, moving the files solved the problem because the data directory does not appear to be used by the program.  Will continue to monitor status.

10/02/03

PAD

Condensation inside of dome.  Removed dome and cleaned condensation from inside.  Reinstalled dome.  RTS at 1608 GMT.

10/1/03

KC

Laptop is not communicating with the instrument. Rebooted laptop at 1456 GMT. RTS at 1511 GMT. Returned after PM at 1600 GMT, laptop not communicating again. Shut down and restarted laptop at 1601 GMT. RTS at 1615 GMT. 

9/29/03

PAD

Computer program locked up on 9/28/03 at 0825 GMT.  Rebooted computer on 9/29/03 at 1408 GMT.  Program came up and is graphing normal.  RTS at 1423 GMT.

9/22/03

KC

There was condensation present in the Dome. Removed and cleaned Dome at 1540 GMT. RTS at 1548 GMT.

9/19/03

KC

Laptop is not communicating with the instrument. Rebooted laptop at 1915 GMT. RTS at 1925 GMT.

9/10/03

KC

Mentor asked that Desiccant be checked and replaced if needed. Replaced Desiccant at 2026 GMT. RTS at 2033 GMT.

9/10/03

PAD

Computer clock was 1 hour ahead of GMT time.  It showed 1557 at 1457 GMT.  Contacted instrument mentor to make sure computer clock was to be on GMT time.  Computer was upgraded recently and set to different time zone. Changed time zone setting and reset computer clock at 2010 GMT.  RTS at 2016 GMT.

9/2/03

KC

Condensation was found inside dome. Removed, cleaned, and resealed Dome at 1503 GMT. RTS at 1515 GMT

8/27/03

KC

SDS is not communicating with instrument. Rebooted SWS desktop at 1505 GMT. RTS at 1515 GMT.

8/22/03

MR

The SWS computer has arrived and needs to be installed.  Reinstalled the computer.  Powered up the computer and the instrument.  The computer started graphing and taking data at 1905 GMT.  RTS at 1915 GMT.

7/24/03

JEM

The SWS computer program was observed locked up.  The computer was totally unresponsive and had to be restarted using the reset button @1423 GMT.  At 1426 GMT, the SWS was back online and operating normally.  RTS @1437 GMT.

7/22/03

MR

The dome has condensation inside.  Removed dome and dried the condensation out.  Replaced dome.  RTS at 1440 GMT.

7/1/03

MR

The dome has condensation inside.  Removed the dome and cleaned the condensation out.  Replaced the dome.  RTS at 1845.

6/30/03

JEM

The SWS clock is 1 minute 30 seconds slow.  Reset SWS clock to correct GMT time @14:47:40 GMT.  RTS @1501 GMT.

6/25/03

JEM

During morning rounds, condensation was observed inside the dome and a pool of water was found inside the trailer on the floor.  Removed the dome from the tube and found a few high spots around the screw holes.  This is most likely the source of the leak.  Wiped the condensation from the inside of the dome.  Ground down the high spots so that the dome could make better contact with the top of the tube.  Replaced the dome and tightened down the screws ensuring that the dome was tight but not in a bind.  If leaks continue a small bead of silicone will be applied around the edges.  RTS @1905 GMT.

6/18/03

JEM

The SDS has not successfully collected any data from the SWS since 5/17/03 @1502 GMT.  The SWS program was found locked up displaying a date and time of 6/17/03 @144759 GMT.  Power cycled the SWS computer @1421 GMT.  The Computer was back online @1424 GMT and the SWS program was operating normally at that time.  RTS @1435 GMT.

6/16/03

JEM

The SWS dome is full of condensation.  Removed the SWS dome @1440 GMT and cleared the condensation from inside.  Checked the desiccant and it was all still blue in color.  Applied a small amount of grease to the O-Ring and reinstalled the dome @1445 GMT.  RTS @1448 GMT.

5/6/03

JEM

The SWS computer was found locked up again.  Power cycled the SWS computer @1605 GMT.  At 1607 GMT the SWS computer was back online and operating normally.  At 1610 GMT, the SWS program unexpectedly froze again.  Several more attempts were made to start the SWS program, but each time the SWS program locked up within a few minutes of being started.  Contacted instrument mentor.  No further action taken at this time.  RTS @1635 GMT.

5/5/03

JEM

The SWS computer was observed locked up.  Restarted the SWS computer @1556 GMT.  At 1602 GMT, the SWS was observed online and operating normally.  RTS @1605 GMT.

4/11/03

JEM

The SWS clock is approximately 2 minutes slow.  Adjusted the clock from 1424 GMT to 1426 GMT.  RTS @1440 GMT.

4/10/03

JEM

The SWS computer was found locked up displaying a time of 1548 GMT.  Closed the SWS program using CTRL-ALT-DEL @1428 GMT.  Restarted the SWS program @1429 GMT.  The SWS was returned to normal operation @1430 GMT.  RTS @1443 GMT.

3/4/03

JEM

The SDS has collected no data from the SWS since 1702 GMT.  Upon checking the SWS, the program was found to be locked up.  Closed the program using CTRL-ALT-DEL  @2130 GMT.   Started the SWS program @2131 GMT.  At 2135 GMT the SWS was observed operating normally.  RTS @2145 GMT.

2/4/03

JEM

The entire Central Facility has experienced a power outage at approximately 1845 GMT.  Power  was restored to the Central Facility at approximately 1945 GMT.  RTS @2000 GMT. 

1/13/03

JEM

The SWS has been red on the Hands page since 01/11/03 @1602 GMT. The SWS program locked up @1520 GMT on 1/11/03.  Rebooted the SWS computer @1607 GMT.  At 1611 GMT, the SWS was observed operating normally.  RTS @1620 GMT.

1/9/03

JEM

The SWS program has locked up.  Closed the program  @1516 GMT using CTRL-ALT-DEL.  Rebooted the SWS computer @1517 GMT.  At 1522 GMT the SWS computer was observed operating normally.  RTS @1529 GMT.

12/31/02

JEM

The SWS program has performed and illegal operation and the computer needs to be restarted.  Restarted the SWS computer @1545 GMT.  At 1555 GMT, the SWS program was running normally.  RTS @1558 GMT.

12/23/02

JEM

The SWS program has locked up.  Rebooted the SWS computer@1505 GMT.  At 1510 the SWS was back online and appeared to be functioning normally.  RTS @1520 GMT.

12/16/02

JEM

Condensation was observed inside the SWS dome.  Removed the dome @1710 GMT so the condensation could be removed.  While drying the dome, the dome popped off of the plate it is mounted on.  Cleaned off the old JB Weld that held the dome on and reapplied fresh JB Weld.  Allowed the JB Weld to set up for a couple of hours inside the trailer where it is warm.  Reinstalled the SWS dome @1950 GMT.  RTS @1955 GMT.

12/12/02

JEM

After being restarted on 12/11/02 @2054 GMT, the SWS software locked up again @2142 GMT.  Used CTRL-ALT-DEL to end the SWS program.  Powered the SWS computer off @1442 GMT.  Powered the computer back on @1443 GMT and at 1445 GMT, the SWS computer was back online.  At 1450 GMT the SWS computer was operating normally.  RTS @1455 GMT.

12/11/02

CAM

SDS instrument status page indicated that no files had been received from the SWS since 1602 GMT.  Checked the instrument and found the Fr_Auto program had locked up at 1532 GMT.  Closed FR_Auto program and restarted at 2054 GMT.  RTS status O.K.

12/6/02

SS

The instrument was not sending files since 1600 GMT.  The instrument computer was noticed to be locked up and was  rebooted at 2045 GMT.  When the reboot was complete, the computer and programs were operating properly.  RTS 2100 GMT.  Status was OK at this time.

11/20/02

MR

The computer was locked up.  Rebooted the computer at 1537 GMT.  RTS at 1552 GMT.

11/18/02

JEM

The SWS program locked up @00:25:04 GMT.  Used CTRL-ALT-DEL to end FR_AUTO.  Restarted FR_AUTO @1520 GMT.  The program was observed running normally @1521 GMT.  RTS @1533 GMT.

11/7/02

KT

There was condensation inside the dome.  At 1933 GMT removed the housing and wiped out the condensation.  At 1941 GMT placed the housing back into position.  RTS at 1948 GMT.

10/22/02

MR

The computer is locked up and is not scrolling data.  Rebooted the computer at 1453 GMT.  Status ok.  RTS at 1510 GMT.

10/16/02

MR

The SWS has locked up.  Rebooted the computer at 2100 GMT.

10/15/02

MR

The program was locked up.  Rebooted the computer at 1455 GMT.  RTS at 1510 GMT status OK.

10/10/02

JEM

The SWS dome has condensation inside it.  Removed the dome @2010 GMT and cleaned the inside.  Replaced the dome ensuring a tight seal @2015 GMT. RTS @ 2020 GMT.

10/3/02

JEM

FR Auto 1.02, has locked up at 14:40:23 on Oct 2nd.  Restarted the SWS computer @1413 GMT.  At 1420 GMT the SWS was operating normally. RTS @1426 GMT.

9/30/02

JEM

FR Auto 1.02 locked up at 14:12:10 GMT on Sept. 30th.  Used CTRL-ALT-DEL to end the FR Auto 1.02 program.  Restarted FR Auto 1.02 @1500 GMT.  System status normal @1510 GMT.  RTS @1515 GMT.

9/19/02

JEM

FR Auto 1.02 has frozen at 00:56:36 GMT.  Used CTRL-ALT-DEL to close FR Auto 1.02. @1433 GMT.  Restarted FR Auto 1.02 @1435 GMT.  Status OK.  RTS @1446 GMT.

9/16/02

JEM

The SWS computer time is off by 1 minute 45 seconds.  Set the clock @15:09:45.  Status normal.  RTS @1520 GMT.

8/19/02

JEM

The computer display for the SWS was observed locked up showing a time of 1435 GMT.  Rebooted the SWS computer @1455 GMT.  Status returned to normal @1459 GMT.  RTS @1500 GMT.

8/9/02

MR

The computer is locked up.  Rebooted the computer at 1433 GMT.  RTS at 1448 GMT.

8/5/02

JEM

The SWS program has frozen @15:14:19.    Rebooted the computer @1429 GMT.  The program started and returned to normal @1437 GMT.  RTS @1444 GMT.

8/2/02

CT

The computer program was terminated. The computer was rebooted at 1445 GMT. Status OK. RTS at 1500 GMT.

7/12/02

KT

The computer was found to have locked up.  The graph and data was not scrolling.  At 1648 GMT rebooted the computer.  At 1653 GMT the computer came on line and the graph appeared and was scrolling.  The values appeared and were scrolling.  RTS at 1703 GMT. 

7/9/02

CT

The computer clock was 2 minutes slow. The computer clock was adjusted at 1411 GMT. RTS at 1426 GMT.

6/28/02

CT

The program was hung up. Rebooted the computer at 1600 GMT. Status OK. RTS at 1615 GMT.

6/27/02

CAM

The SWS Fr_Auto program was locked up.  The last update was at 1856 GMT on 6/25.  Network Administrator checked for data files on SDS and found that does collect a data file every hour when FTP is running and then deletes it from the SWS computer.  The last file collected was at 1800 GMT on 6/25.  Restarted the program at 1940 GMT.  Instrument mentor modified the asd.cfg file to force the system to optimize every hour.  The system optimized and the spectrum amplitude increased and now appears normal.  RTS status O.K. at 1955 GMT.

6/25/02

MR

The instrument mentor requests that the WSFTP program be shut down so that the instrument program can be tested to see if it creates data files.  Shut the program down at 1353 GMT.  Allowed the instrument to run for 5 hours and observed that the program did create 6 data files.  Notified the mentor and he requested that the FTP program be restarted.  Before restarting the program I placed a copy of the data files in the C:\CHRIS folder.  I then restarted the WSFTP program.  The mentor said he would look at the script program and let us know what needs to be done.  RTS at 1915 GMT.

6/24/02

CAM

The programmer from ASD (Terry) called to T/S the SWS problem.  The instrument does not appear to be writing data files correctly since it was returned from ASD.  The SWS program was stopped at 1747 GMT.  Emailed a copy of the current log (.olg) file to ASD programmer for diagnostics.  Restarted the program and RTS at 1812 GMT.  T/S to continue.

6/21/02

CAM

The SWS program was hung up and not updating.  The last update was at 2054 GMT on 6/20.  Attempted to exit the program using Alt-Q but it would not respond.  Minimized the program and closed the window.  Restarted the program from the desktop shortcut and returned operational at 1520 GMT.  Notified mentor and vendor of the problem.

6/20/02

CAM

The SWS software needed to be modified to run the Fr_Auto program rather than the Fr program and use the #7 calibration files.  Modifed the Fr_Auto shortcut to run the #7 calibration files.  Added the shortcut to the startup group so that the program will start automatically on bootup.  Copied 1i682_7.raw, Ni682_7.raw, Lmp682_7.ill and Bse682_7.ref to c:\auto\source.  Rebooted the computer and verified the correct program started.  RTS at 2041 GMT.

6/14/02

MR

The system lost power on 6/13/02 due to power outage to the site.  The instrument power switch was turned off until power was restored to the site.  Power has been restored to the site.  Powered on the instrument at 1400 GMT.  System came back up and started taking data at 1403 GMT.  RTS at 1415 GMT.

6/13/02

MR

The site has lost power due to storms in the area at 1415 GMT.  The instrument is shut down at this time.

6/11/02

MR

The *SWS has returned from the factory and needs to be reinstalled.  Reinstalled the computer system, dome and instrument. Powered on the instrument at 1630 GMT. The computer came up and started the FR auto program but immediately stopped. Clicked on the FR Auto icon and the program restarted. Notified the mentor.  *New – 682 (cal #7)

4/23/02

MR

The instrument mentor wants the *SWS instrument and computer removed and sent back to ASD for repairs. Removed the dome and instrument.  Installed the plug plate in the enclosure.  Removed the computer but left the keyboard, mouse and monitor.  Packed the instrument up in the shipping container and returned it to shipping and receiving for shipment.  *Old - 682

3/26/02

Mentor

Mentor on site to make modifications to system and T/S the software problems.  Mentor modified the fiber optic so that it now points directly at the sky rather than measuring irradiance by pointing at the spectralon panel.  The dome was also replaced with a much smaller version.  Mentor worked with ASD and determined there is a problem with the instrument, not the software.  Mentor will further discuss this with ASD and advise Site Ops if the instrument needs to be returned to the vendor.  System remains off line at this time.

3/4/02

CT

Due to power outage the instrument was found off. Rebooted the computer at 1527 GMT. RTS at 1542 GMT.

2/19/02

CT

The graph was not producing data results. Power cycled the computer at 1701 GMT. This did not fix the graphing problem. The mentor has been contacted about the problem. No further CM at this time RTS at 1716 GMT.

2/11/02

MR

The AC power to the site was lost on 1/30/02 due to local ice storm.  Power to the site was found to have been restored.  Powered up the instrument at 1530 GMT.  Status OK.

1/28/02

CT

The graph display was blank. Rebooted computer at 1702 GMT. This did not correct the problem. At 1703 GMT the computer was power cycled. This did correct the problem. RTS at 1717 GMT.

1/25/02

CT

The display screen was found blank at 1536 GMT. Attempted to correct problem with a soft reboot at 1536 GMT. This did not correct the problem. At 1540 GMT the computer was power cycled. The computer came online at 1542 GMT. RTS at 1551 GMT.

1/24/02

CT

The computer screen was blank. The computer was power cycled  at 1518 GMT. RTS at 1533 GMT.

1/17/02

MR

The SWS has been returned from calibration and needs to be installed.  Installed the *SWS.  The cracked dome was returned and used during installation.  There is also a new smaller dome in the shipping crate but no mounting hardware.  The mentor stated he would have to make a trip here to set up the new dome.  Started the computer at 1627.  At 1631 I noticed the computer would reboot itself if the AUTO_FR screen was maximized but the computer would run correctly if the window was minimized.  Notified the mentor of the problem and he said he would contact the vendor about the problem and said to just run the program minimized for now.  RTS the unit at 1645 GMT.  *New - 682

10/26/01

CJ

Found the computer program locked up at 1436 GMT. Rebooted the computer. RTS at 1451 GMT.

10/25/01

CJ

Program indicated that the graph needed to be optimized.  Optimized the graph at 1520 GMT. RTS at 1535.

10/16/01

CJ

The graph needed to be optimized.  At 1519 GMT the graph was optimized.  RTS at 1533 GMT. 

10/5/01

CAM

Instrument mentor requested that the ‘Save Files’ option be restarted.  Set the time interval to 1 minute and started the save option at 1600 GMT.

9/17/01

MR

The SWS had no graph displayed.  Rebooted the computer at 1426 GMT.  System came back up and started working at 1428 GMT.  Restarted the data saving at 1428 GMT.

9/7/01

CAM

Instrument mentor requested that the sample interval be changed from 30 seconds to 1 minute.  Made the requested change at 2050 GMT.

8/31/01

CAM

The SWS computer was not saving files.  The computer was rebooted on 8/30 and the save settings were lost.  Optimized the instrument using the control pull down menu functions at 1935 GMT.  Adjusted the number of spectrum value in the configuration menu to 290.  Set the spectrum save path to c:\auto\readfrom.  Set the maximum number of files to save as 9999 and the sample interval to 00:00:30.  RTS at 1940 GMT.

8/30/01

MR

The SWS was not reporting any data. T/S found that the system was running.  Rebooted the computer at 1443 GMT.  RTS at 1458.

8/27/01

CAM

Instrument mentor requested that the SWS be optimized.  Optimized the instrument using the control pull down menu functions at 1730 GMT.  Set  the sample interval to 00:01:00.  RTS at 1740 GMT.

8/21/01

CAM

Instrument mentor requested that the SWS be optimized.  Optimized the instrument using the control pull down menu functions at 1640 GMT.  Adjusted the number of samples value in the configuration menu to 290.  Set the spectrum save path to c:\auto\readfrom.  Set the maximum number of files to save as 9999 and the sample interval to 00:00:30.  RTS at 1655 GMT.

8/20/01

MR

The power was lost to the site on 8/15/01 from 0030 GMT to 0400 GMT.  The computer was found off. Powered on the computer at 1410 GMT. System came up and started running at 1413 GMT.  RTS at 1425 GMT.

8/8/01

CAM

The SWS graph that was being displayed did not appear normal and was not updating.  Shut the program down and restarted it at 1512 GMT with no change in status.  Rebooted the computer at 1520 GMT but there was still no change.  Contacted the instrument mentor and optimized the instrument settings at his request using the pull down menu functions.  Started the save files option and system RTS status O.K. at 1620 GMT.

7/25/01 – 7/26/01

Mentor

Instrument mentor Peter Pilewskie and Larry Pezzolo on site to T/S the SWS and get familiar with the instrument.  Found that the instrument would not operate properly with the specialized Fr_Auto program but would run with the regular Fr program.  Mentor downloaded the latest version of the Fr program from ASD’s website and system is now operating under this new software.  Several issues must be resolved before the SDS is able to accept the new data format and the mentor will take the necessary steps.  In the meantime, the data is being written to the readfrom directory on the PC and the mentor will pull it off via FTP.  The dome was inadvertently cracked during maintenance and will need to be replaced.  The crack was temporarily sealed with silicone until the mentor can ship a new dome.  System RTS at 1800 GMT on 7/26.

7/10/01

CAM

The new SWS instrument mentor (Peter Pilewskie) contacted Site Ops to request some information about the status of the SWS.  Connected the SWS to the computer and powered it on at 1830 GMT.  The program started normally and the computer and instrument were communicating.   However, the displayed data did not appear reasonable.  There was a large spike around 1000 nm.   The mentor suspects this could be a problem with the calibration.  Used the Print Scrn option to capture an image of the spectrum and emailed it to the mentor for evaluation.  System left running at this time.

8/8/00

MR

Shut the instrument down at 1920 GMT as per mentors request until further notice.

8/3/00

KT

At 1412 GMT the instrument computer was found locked up. At 1414 GMT rebooted the computer. At 1416 GMT the computer came on line. RTS at 1429 GMT.

8/2/00

KT

At 1410 GMT the computer was found locked up. At 1411 GMT rebooted the instrument computer. At 1413 GMT the computer came on line and the data started to scroll. RTS at 1425 GMT.

8/1/00

KT

At 1406 GMT observed that the instrument computer was locked up. At 1407 GMT rebooted computer. At 1409 GMT computer came on line and data started scrolling. RTS at 1421 GMT.

7/27/00

KT

At 1705 GMT the instrument computer was found to be locked up. At 1706 GMT rebooted computer. At 1708 GMT the computer came on line. RTS at 1726 GMT.

7/21/00

KT

At 1616 GMT the instrument computer was found locked up. At 1617 GMT rebooted the computer. At 1622 GMT the computer came on line and the data started to scroll. RTS at 1631 GMT.

7/20/00

KT

There was a power outage that occurred at the CF. The power was restored at 1427 GMT. Instrument cables were reseated. Instrument came back on line. RTS at 1530 GMT.

7/3/00

MR

Collections have failed with the SWS. The SWS computer seemed to be working OK. The system was collecting data and storing it in the ASD file. Rebooted the computer. Later checks on the system found the collections were restored. Also loaded the new calibration files.

6/30/00

MR

Reinstalled the *SWS instrument. Powered up the instrument at 2030 GMT and enabled the collections and ingest on the site ops page. ASD made repairs, installed a new fan, dome, and calibrated the instrument. Status O.K. *New - 682

5/30/2000

CAM

SWS SN 682 was received back from Analytical Spectral Devices after being repaired. Upon opening the instrument case, it was found that the heat sinks and fan were loose inside the case. The data and fiber optic cables and instrument were covered with heat sink compound. The screws that should have secured the heat sinks and fan were also loose inside the case. Cleaned up the instrument and cables to remove the excess heat sink compound. Noticed a spot on the fiber optic cable that looked like it had been pinched but was unable to determine if it was enough to cause damage to the fiber. Attempted to secure the heat sinks using the screws that were loose inside the case. The screws would not tighten the heat sinks adequately enough to hold them securely in place. Used longer screws with nuts to secure the heat sinks in place. Installed the instrument and connected to the computer. Powered on the instrument and then powered on the computer. The Fr Auto program started briefly and then closed without ever displaying a graph. Numerous attempts at restarting the program had the same result. Spoke with Bob Faus at ASD and he suggested returning the instrument for inspection. Suspect the instrument may have been damaged in shipment. Packed *SWS SN 682 to return to ASD. Reference PO# OA-07389 RMA# 053100

2/10/2000

CAM

Jeff Griffin requested that the SWS be pulled and returned to the vendor for calibration and to be checked out. Pulled the *instrument, power supply, dome, and Cosine reflector and shipped to ASD, Inc. *Old - 682

2/7/00

GRS

The SWS was not plotting data. Stopped all programs and booted the computer at 1534 GMT. The instrument and computer did not have a link after the computer reboot. Shut down the computer at 1541 GMT. Power cycled the instrument at 1542 GMT. Powered up the computer at 1543 GMT. The computer started its boot up but rebooted about half way through its boot process. The computers watchdog timer keeps rebooting the computer before it finished its previous reboot. Disabled the watchdog timer but the computer keep rebooting. Pulled the watchdog timer out of the instrument but the computer would stop when it was time for the watchdog timer to kick in. The computer would not reboot but it would stop during its boot up. Called the Mentor and worked with him to remedy the problem. After several attempts to fix the instrument the mentor asked us to copy some files to a disk and email them to him. He wanted to view the programs. He advised us to shut down the instrument and computer until he had time to view the programs. He called us the next day on 2/8/00 and asked us to pull the instrument and ship it back to ASD for evaluation and annual calibrations. Left the instrument down until it could be pulled and sent back to ASD.

1/6/2000

CM

While retiling the Optical trailer floor, contractors plugged a compressor into a UPS outlet which caused the breaker to trip. The UPS and all Optical trailer instruments lost power and had to be restarted. Removed the compressor from the UPS outlet and reset the breaker. Restarted all Optical trailer instruments. RTS status O.K.

12/28/99

MK

Due to power line maintenance starting at 1500 GMT all instruments will need to be shut down. Shutdown SWS at 1435 GMT, restarted SWS at 1618 GMT. RTS at 1618 GMT.

10/7/99

GRS

The data has been reported as being bad by the Mentor and that the plus 5 and 15 voltages were spiking. A check of the instrument showed no visual problems. The 5 and 15 voltages were well within specs. The data looked good. The program, program times, and computer times were all correct. Jeff Griffin believed that since the SWS program was running normally and that the data received on their end was bad that it might be the A to D converter. He suggested that the computer be rebooted. Shut down the SWS program and computer at 2132 GMT. Rebooted the computer. The computer then started rebooting on it's own about every 30 seconds. It was stuck in a loop. Powered off the computer and disconnected the instrument from the computer. The computer booted normally. Stopped the computer and then reconnected the instrument. Booted the computer, but the computer then started rebooting on it's own about every 3 seconds. Powered off the computer. Disconnected the instrument from the computer. Booted the computer. It came up normally. Disabled the watchdog timer. Reconnected the instrument to the computer and then booted the computer. The computer came up normally and started taking data. Reenabled the watchdog timer, stopped the program, and rebooted the computer. This time the computer booted normally and started taking data. For some reason the watchdog timer was stuck in a loop. RTS instrument at 2214 GMT.

9/10/99

MR

Reset the system clock. RTS at 1450.

9/1/99

MR

They system clock was off by more then 10 secs. Reset the system clock to correct GMT time.

8/9/99

MR

Rebooted the computer due to site power construction. RTS at 1530.

8/3/99

MR

System clock was off by 10 secs. Reset the system clock. RTS at 1420.

7/22/99

MR

Reset the system time. Off by 8 sec. RTS at 1407.

3/24/99

CAM

SDS has not collected files from the SWS since 1939 GMT. SDS comlog indicates 'FTP Problem Occurred During Collection Process'. Rebooted the SWS computer at 2142 GMT due to network cabling upgrades being done in the optical trailer. RTS at 2144 GMT. SDS collections then returned to normal.

3/15/99

MW

The CF lost power due to a snow storm on 3/14/99 at 0030 GMT. The SWS has been down since that time. Powered the system back on after AC power was restored. RTS at 2035 GMT. Status OK.

3/9/99

MW

The SWS came back yesterday at 2215 GMT from diagnostics and repair at ADS. Installed the *SWS 1445-1519 GMT. Powered the instrument on before powering the computer on as per mentor direction. RTS at 1519 GMT. *New - 682. Calibration #3.

3/8/99

CAM

Installed new asd.cfg, default.cfg, asd.ini, cos682_3.raw, and lmp682_3.ill files to c:\auto\source directory in preparation for receipt of the repaired SWS. Updated all Fr_auto shortcuts to reflect the new calibration number (c:\auto\source\Fr_auto 682 3). Instrument should arrive late today. The vendor found that a linking to one of the drive motors failed and was repaired under warranty.

3/1/99

CAM

Dave Beal sent a copy of Wavecal.exe and Egavga.bgi and requested that they be copied to c:\auto\source and Wavecal be executed. Copied the files and executed Wavecal 682 2 at 2210 GMT. The results were random with SWIR Size varying widely from 592 up to 800. 592 was the most common result but Dave advised that the value varied far more than it should. Similarly, the index channel reading varied from 26 to 377 with 26 being the most common and occurring when SWIR Size was 592. Dave suggested that the instrument be removed and sent to ASD for diagnostics/repair. Powered the system down at 2225 GMT. Pulled the *instrument, power supply, dome, and Cosine reflector and shipped to ASD, Inc. on 3/2/99. System off line until returned from vendor. *Old - 682. Calibration #2.

3/1/99

CAM

Dave Beal sent a new asd.cfg file to optimize every half hour. Loaded the new file and restarted the program at 1700 GMT. The new file produced no change in status. The computer continues to reboot itself.

2/23/99 - 2/26/99

CAM

The SWS computer appears to be having problems communicating with the instrument. As a result the watchdog timer is continuously rebooting the computer. Upon initial boot up the system comes up normal with voltages displayed and Dark Current Collections showing 50 samples in progress. However no graph is displayed and as soon as the Spectrum Average shows 400 samples in progress the system reboots. The process is repeated continuously. Working with Dave Beal, ASD Product Specialist, the following maintenance was performed. On 2/23/99: Sent copies of ASD.CFG, ASD.INI, and 90540320.OLG to Dave for evaluation. Dave suspected that the optimization parameters needed to be changed and sent a new ASD.CFG file. Copied the new file to c:\auto\writeto and c:\auto\source directories. Restarted the SWS software at 2004 GMT but there was no change in status. Dave later sent test software to see if the instrument was working at all. Installed the test software to c: and ran Fr.exe program from 2140 to 2150 GMT. The test program worked and the instrument itself is operational. Sent copies of Spectrum.000 and config.sys to Dave for evaluation. On 2/25/99: Dave sent new config.sys and ASD.CFG files. Copied the new ASD.CFG to c:\auto\writeto and c:\auto\source directories. Restarted SWS software at 1655 GMT. This time the program ran normally with the graph and all data displayed. Did not load the new config.sys file at Dave's request. Stopped the program at 1705 GMT and ran ScanDisk and defragged drive c:. ScanDisk found no errors on the drive and it was only 3% fragmented. Rebooted the computer at 1858 GMT and RTS status O.K. at 1859 GMT. At 1940 GMT the computer rebooted itself but came back up O.K. Contacted Dave and he suggested reloading the executable program Fr_Auto.exe plus numerous miscellaneous files. Reloaded software and rebooted computer at 2000 GMT. The program started O.K. but the time was 1 hour slow. Corrected the time and RTS at 2010 GMT. At that time the system was running normally. However status check the following day at 1500 GMT found the system was once again rebooting continuously just as it had been initially. Suspect a problem remains with the software. This is a unique system that only Dave has the knowledge of to provide adequate technical support and he is out of town until 3/1/99. Closed the SWS program to prevent the continuous reboots at 1700 GMT. System is off line until further support is available from ASD.

2/23/99

CAM

The SWS collections began failing again at 2001 GMT on 2/22. SDS indicated there were 'No Files to Collect'. Check of instrument found once again there was no graph displayed on the monitor. The temperature and voltage readings are being reported, but the watchdog timer continues to reboot the computer. Stopped the Fr_Auto program at 1452 GMT on 2/23. Copied the ASD.INI and ASD.CFG files and sent to Dave Beal at Analytical Spectral Devices. Dave will go over the files to see if changes to the optimization process are necessary. System off line at this time.

2/22/99

KT

The SWS reported its last successful collection at 1843 GMT on 2/19/99. After that time collection of data ceased. At 1558 GMT notified Chris Martin that the screen was in windows mode. After instruction the FR Auto switch was clicked and the graph screen appeared. Notified Chris Martin that graph was on and data collection was in progress. The SWS was RTS at 1559 GMT on 2/22/99

2/19/99

CM

Daily status check found that the watchdog timer was continuously rebooting the SWS computer. The program would come up momentarily with no graphical data displayed and then the computer would reboot. Verified that the instrument had power and was running. Shut the computer completely down and restarted at 1440 GMT. The computer came up in the same state. The temperature and voltages are being displayed on the monitor so the PC is at least partially communicating with the instrument. However the Spectrum Average display reads 'Samples: Unknown None' for approximately 1 minute following reboot, then the Spectrum Average display reads 'Sample: 400 In Progress' momentarily and the computer reboots. The cycle is repeated over and over. At no time does the graph appear on the monitor. Notified mentor via voice mail and left the system powered on as is.

2/16/99

MR

Reset the time on the computer. It was off by 00:01:25.

11/20/98

SS

The instrument is not transferring files. Rebooted the computer at 1735 GMT.

9/29/98

KT

At 1833 GMT per mentor instruction, the SWS computer was powered off and restarted. At 1840 GMT the screen came on and the data appeared to be scrolling like normal. RTS at 1843 GMT.

6/30/98

CAM

Received repaired SWS from vendor. Installed SWS from 1330 to 1350 GMT. Powered system on at 1353 GMT. The asd.cfg configuration file had been modified prior to removing the instrument as part of T/S. Restored asd.cfg file back to original format. Restarted software and RTS status O.K. at 1430 GMT. *New - 682

6/16/98

CAM

Removed *SWS from service at 1330 GMT. Will return to vendor for repair. *Old - 682

6/15/98

CAM

Received phone call from Dave Beal from Analytical Spectral Devices, SWS vendor. Provided phone support and emailed copies of SWS cfg and olg files to vendor for them to analyze. Vendor advised the A/D converter or internal power supply has likely failed. Powered system completely down at 2110 GMT. Will pull system and return to vendor for warranty repair when proper tool is obtained for removal.

6/11/98

CAM

The site experienced a power bump at 2333 GMT on 6/10 and the following SWS collection (2336 GMT) failed with 'No Files To Collect'. All of the ensuing collections failed as well with the same SDS Comlog error message. Instrument status check found that the FR_Auto program was running but there was no graphical representation of the data displayed. The Spectrum Average status box indicated 'Samples: Unk' and 'None'. Rebooted computer at 1350 GMT but the display came up the same again (no graph). Checked the data directory (c:\auto\source\inprog) and there were no files there. Checked the c:\auto\readfrom directory and there was a file there. Following the reboot the SWS collection was successful at 1436 GMT but the file it collected was 81151200.ASD which is from April 25. The 1536 GMT collection failed again with 'No Files To Collect'. At 1600 GMT there was still no graph displayed on the monitor. Power cycled instrument at 1610 GMT and rebooted computer but no change in status. Left message with mentor and T/S in progress.

5/15/98

SS

Due to power failure restarted instrument at 1525 GMT.

5/7/98

SS

Found the south west guy wire undone and lying next to port hole. Reset guy wire and tightened. RTS at 1345 GMT.

4/8/98

TG

Instrument mentor reports that he can not delete 2 files from the ftd directory. T/S of system indicates the wftpd.exe software had locks on the 2 files when it tried to transfer the two files. Shutdown wftpd.exe and removed the two files, 80750646.ASD and 80931000.ASD. Restarted wftpd.exe. Checked user permissions on C:\AUTO directories to allow Instrument Mentor to delete/rename files.

3/31/98

ZW

The SWS instrument computer time was reading 1610 at 1510 GMT. Set the instrument computer time from 1610 to 1510 GMT. Then set the program time to match the computer time. Status ok. RTS 1525 GMT.

3/6/98

TG

Mentor requested that the computer be connected to development network. Configured computer software for network address of 198.124.98.53 and rebooted computer at 1621 GMT. RTS system at 1624 GMT.

2/26/98

Mentor

System installed by mentor (Jeff Griffin). Spectrometer *SN 682. Calibration #2.